OpenStudy (anonymous):

simplify 3xy-6y/4x^2y-8xy

OpenStudy (anonymous):

3xy-6y/4x^2y-8xy =3y(x-2)/4xy(x-2) eliminate (x-2) =3y/4xy =3/4x